OpenStudy (anonymous):

which ordered pair is a solution of the equation y=x-4

OpenStudy (anonymous):

(2,6) (6,2) (-2,6) (-6,2)

OpenStudy (anonymous):

(x, y) also, the eq states y to always be 4 smaller than x

OpenStudy (anonymous):

wait, what?

OpenStudy (phi):

a solution means when you replace the x and y with the numbers they give you, the equation is true. for example, is (2,6) a solution of y = x -4 ? (2,6) is short for let x=2 and y=6. use those numbers in y=x-4 6 =? 2 -4 2-4 is -2 and we get 6 = ? -2 NO. (2,6) is not a solution test the other pairs of numbers. Do any work ?